The perimeter of a rectangular deck is 228 feet. It is 69 feet long. How wide is it?
Answer by TimothyLamb(4379)

---
p = 2(L + w)
228 = 2(69 + w)
228 = 138 + 2w
2w = 90
w = 45 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php